Find the values of each of the following polynomials for the given values of the variable.
-x^2 - 5x - 6, x = -3 and x = -2

Answer by homeworkhelpanytime(21)

-x^2 - 5x - 6, x = -3 and x = -2
when x = -3
= -(-3)(-3) - 5*(-3) - 6
= -9 + 15 - 6
= 0
when x = -2
= -(-2)(-2) - 5*(-2) - 6
= -4 + 10 - 6
= 0